# CrosGrpsABS: Cross-Attention over Syntactic and Semantic Graphs for Aspect-Based Sentiment Analysis in a Low-Resource Language

### v1

@@ Abstract
Index Terms:
Aspect-based Sentiment Analysis, Syntactic Graph, Semantic Graph, Hybrid Graph Neural Network, Low-Resource, Cross-Attention
Impact Statement– The CrosGrpsABS framework advances aspect-based sentiment analysis by overcoming the limitations of isolated syntactic or semantic modeling. Integrating syntactic dependency and semantic similarity graphs through bidirectional cross-attention, it dynamically aligns structural and contextual cues for deeper sentiment understanding. This unified graph-based approach delivers state-of-the-art performance across both Bengali (a low-resource language) and benchmark English datasets. By enabling interpretable cross-lingual adaptability and fine-grained aspect detection, CrosGrpsABS not only enhances ABSA accuracy but also lays the foundation for robust, language-agnostic sentiment systems in diverse real-world applications such as opinion mining, social media analytics, and resource-constrained NLP.
I Introduction
@@
Recent advances leveraging transformer-based models such as BERT [8], RoBERTa [9], and graph-based neural network architectures [2, 10] have significantly propelled ABSA research, particularly for high-resource languages such as English. However, other studies indicate that these models rely heavily on substantial annotated data and language-specific resources, creating significant barriers for low-resource language applications [10, 11]. These challenges lead to several key research questions (RQs): (RQ1) How do syntactic dependency and semantic similarity graphs, individually and in combination, affect the performance and interpretability of aspect-based sentiment analysis models? (RQ2) How can existing transformer-based and graph-based ABSA techniques be effectively adapted across languages with varying resource availability, such as low-resource languages like Bengali and high-resource languages like English?
To address the research questions, we propose a novel framework, CrosGrpsABS, that unifies syntactic and semantic graph representations through a bidirectional cross-attention mechanism. In response to RQ1, the framework investigates how syntactic dependency and semantic similarity graphs, individually and jointly, contribute to both the performance and interpretability of ABSA models. For RQ2, it explores how such a hybrid model can be effectively adapted across languages with varying resource availability, particularly focusing on Bengali as a low-resource language and English as a high-resource benchmark. CrosGrpsABS constructs two complementary graphs: a syntactic graph that captures local grammatical relationships using dependency parsing tailored to the linguistic characteristics of Bengali, and a semantic graph that captures global contextual relationships by computing cosine similarities between transformer-based word embeddings. These graphs are dynamically fused through a bidirectional cross-attention mechanism that aligns and refines node representations based on cross-graph signals, allowing syntactic and semantic cues to reinforce each other. This design not only enhances sentiment prediction but also offers interpretability by highlighting influential syntactic structures and semantic associations.
@@
  • ARQ1: We propose a unified graph construction that encodes syntactic dependencies and semantic similarities via transformer embeddings, then integrates them with bidirectional cross‐attention to enhance both interpretability and performance in aspect‐based sentiment analysis.
@@
  • ARQ2: We introduce CrosGrpsABS, a hybrid ABSA framework that combines transformer-based and graph-based techniques. Experiments on low-resource Bengali and high-resource English datasets (SemEval 2014 Task 4: Restaurant and Laptop) show its strong generalization and effectiveness across languages with varying resource levels.
@@
The remainder of this paper is organized as follows. Section II reviews the related work on ABSA, graph-based models, and sentiment analysis for low-resource languages. Section III presents the architecture of the proposed CrosGrpsABS. Section IV describes the experimental setup, and discusses the results and compares them with existing baseline models. Finally, Section V concludes the study and suggests future research directions.
II Literature Review
@@
Aspect-Based Sentiment Analysis (ABSA) has witnessed substantial progress in high-resource languages, particularly English, owing to the availability of large-scale annotated datasets, advanced language tools, and powerful pre-trained language models. Transformer-based architectures such as BERT [12], RoBERTa [13], and their specialized ABSA variants have significantly improved sentiment classification performance by capturing complex contextual dependencies through self-attention mechanisms [14]. These models often serve as the backbone for sentence-level and aspect-level classification tasks. Beyond sequential models, graph-based approaches have gained prominence for their ability to incorporate syntactic and semantic structures explicitly [7]. Dependency tree-based models [15, 16] and graph neural networks (GNNs) such as GCNs and TextGCNs [17, 18] have been employed to construct graph representations of sentence structure, enabling the learning of relational patterns between aspect and opinion terms. Semantic graphs based on similarity measures [19] further enhance global contextual understanding [20, 21]. However, many existing models in high-resource contexts treat syntactic and semantic features independently, limiting their capacity to leverage their complementary strengths for interpretability and performance.
@@
Despite the global prevalence of languages like Bengali, ABSA research in low-resource settings remains underdeveloped. Bengali-specific sentiment analysis efforts have predominantly addressed coarse-grained tasks at the document or sentence level [6, 22], with limited exploration of fine-grained ABSA. The scarcity of annotated aspect-level datasets, lack of robust dependency parsers, and the unavailability of optimized Bengali-specific pre-trained models constrain the development of syntactic and semantic representations [23]. Furthermore, Bengali’s complex morphological structures and syntactic flexibility introduce challenges not observed in high-resource languages. Recent studies have attempted to adapt multilingual transformers (e.g., mBERT, XLM-R) for Bengali ABSA [24, 25], but performance often suffers due to mismatches in linguistic structure and domain specificity. Graph-based methods are even less explored in Bengali due to the difficulty of building accurate syntactic trees and semantic graphs without language-specific tools. This underscores the need for an adaptable ABSA framework that can overcome these limitations by integrating syntactic and semantic structures.
Although syntactic and semantic cues are crucial for ABSA, most models treat them separately. Recent works have tried to bridge this gap with attention or fusion mechanisms, but few offer a unified, interpretable integration. Many models fail to capture cross-level interactions in a structured, bidirectional way and lack explainability, especially in low-resource settings like Bengali. Moreover, cross-lingual adaptability is rarely explored through graph-based techniques, limiting multilingual application. This highlights the need for an interpretable ABSA framework that integrates both syntactic structure and semantic context while being adaptable across languages. To address this, we propose CrosGrpsABS, which unifies syntactic and semantic graphs using a novel bidirectional cross-attention mechanism. Our model effectively tackles RQ1 (graph interaction and interpretability) and RQ2 (cross-lingual adaptability), showing promising results in both Bengali and English ABSA benchmarks.
III Proposed Methodology
[…]
@@ III-A Syntactic Graph
III-B Semantic Similarity Graph
@@
III-C Graph Attention Networks (GAT)
@@
where $\mathbf{W}$ is a learnable weight matrix that projects node features into an attention space, and $\mathbf{a}$ is a learnable attention vector. The concatenation $[\mathbf{W}\mathbf{h}_{i}\,\|\,\mathbf{W}\mathbf{h}_{j}]$ allows the attention mechanism to model interactions between the node and its neighbors, as shown in Equation (3). A LeakyReLU activation function is applied before computing the softmax normalization across all neighbors.
@@
III-D Bidirectional Cross-Attention
@@
III-E Transformer Encoder Refinement
@@
III-F Aspect Representation Extraction
@@
where $|M|$ denotes the number of tokens forming the aspect, and $[1,\ldots,M]$ are the corresponding indices in $\mathbf{H}_{\text{aspect}}$. This pooling step yields a compact embedding $\mathbf{z}_{\text{aspect}}$ that concentrates on the region of the sentence determining sentiment polarity. By focusing on aspect-specific dependencies, the model can disentangle general context from cues most pertinent to the aspect, ultimately enhancing its ability to capture fine-grained sentiment relationships in ABSA.
TABLE I: Data distribution for the four domain-specific ABSA datasets in Bangla.
@@
TABLE II: tatistics of SemEval 2014 Task 4 datasets (Laptop and Restaurant) with sentiment distribution.
@@
TABLE III: Hyperparameter Settings and Additional Techniques
@@
TABLE IV: Comparative performance of three embedding-based approaches (mBERT, RoBERTa, BanglaBERT) on four domain-specific Bengali ABSA datasets. Acc and F1 are reported in percentage. The best results in each column are highlighted in bold.
@@
[…]
@@
TABLE V: Performance comparison of our proposed model with baseline models on multiple datasets. The best results are highlighted in bold, and the second-best results are underlined.
@@
[…]
Roberta-base (Zero-shot) [9] 50.75 46.51 52.53 48.59 63.24 59.07 57.65 51.68
Roberta-large (Zero-shot) [9] 54.52 57.91 68.82 67.46 66.08 65.58 71.82 70.68
Roberta [9] 56.34 50.41 68.28 68.16 74.32 74.46 80.13 79.91
BanglaBERT [32] 56.81 41.16 48.39 31.56 68.24 55.36 54.97 38.99
@@
III-G Highway Gating Mechanism
TABLE VI: Comparison of Different Models on ABSA Tasks (SemEval 2014 Task 4 datasets: Restaurant and Laptop)
@@ III-G Highway Gating Mechanism
TABLE VII: Ablation study results for our proposed model across multiple datasets. The best results are highlighted in bold.
@@
$\bar{\mathbf{h}}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{H}_{\text{refined}}[i],$ (14)
We then define two functions acting on $\mathbf{u}$. The first, $\mathbf{H}(\cdot)$, is a feed-forward transformation (e.g., a ReLU-activated linear layer) that generates a candidate set of new features. The second, $\mathbf{T}(\cdot)$, is a gating function that produces a sigmoid output indicating how much of these new features should be integrated [39]. Specifically, we compute
$\mathbf{z}_{H}=\mathbf{T}(\mathbf{u})\,\odot\,\mathbf{H}(\mathbf{u})+\bigl{(}1-\mathbf{T}(\mathbf{u})\bigr{)}\,\odot\,\mathbf{u},$ (16)
where $\mathbf{T}(\mathbf{u})=\sigma\bigl{(}\mathbf{W}_{T}\mathbf{u}+\mathbf{b}_{T}\bigr{)}$. In Equation (16), the gate $\mathbf{T}(\mathbf{u})$ balances the incorporation of the transformed features $\mathbf{H}(\mathbf{u})$ with the original input $\mathbf{u}$, preserving skip connections to facilitate gradient flow. The resulting $\mathbf{z}_{H}$ provides a harmonized fusion of aspect-aware signals and global context, enhancing the model’s capacity to capture nuanced sentiment relationships. Consequently, $\mathbf{z}_{H}$ serves as the final, enriched representation passed to downstream classification layers.
III-H Classification Layer

### v2

# Conceptualization of this study, Methodology, Data Curation, Software, Writing - Original Draft

@@ Abstract
keywords:
[…]
1 Introduction
@@
Recent advances leveraging transformer-based models such as BERT [8], RoBERTa [9], and graph-based neural network architectures [2, 10] have significantly propelled ABSA research, particularly for high-resource languages such as English. However, other studies indicate that these models rely heavily on substantial annotated data and language-specific resources, creating significant barriers for low-resource language applications [10, 11].
These challenges lead to several key research questions (RQs): (RQ1) How do syntactic dependency and semantic similarity graphs—individually and in combination—affect the performance and interpretability of aspect-based sentiment analysis models? (RQ2) How can existing transformer-based and graph-based ABSA techniques be effectively adapted across languages with varying resource availability, such as low-resource languages like Bengali and high-resource languages like English?
Figure 1: Comparison of syntactic and semantic distances across different aspects (Car, Mobile, Movie, and Restaurant). The results indicate that syntactic distances (blue) are generally larger than semantic distances (red), highlighting the need for a mechanism that effectively integrates syntactic and semantic information for aspect-based sentiment analysis in Bengali texts.
Syntactic distance measures the structural difference between words in a sentence based on grammatical rules, while semantic distance assesses the conceptual similarity between words using their contextual meanings. In the context of aspect-based sentiment analysis, these distances help in understanding how words relate to one another from both a structural and a meaning perspective. Figure 1 shows the average syntactic and semantic distances for four aspects: Car, Mobile, Movie, and Restaurant. Syntactic distances, represented in blue, tend to have higher values indicating more variability in grammatical relationships, whereas semantic distances, shown in red, are generally lower, reflecting more consistent contextual connections. This comparison underscores the importance of integrating both types of information to capture the full spectrum of linguistic nuances in sentiment analysis.
To address the research questions, we propose a novel framework, CrosGrpsABS, that unifies syntactic and semantic graph representations through a bidirectional cross-attention mechanism. In response to RQ1, the framework investigates how syntactic dependency and semantic similarity graphs—individually and jointly—contribute to both the performance and interpretability of ABSA models. For RQ2, it explores how such a hybrid model can be effectively adapted across languages with varying resource availability, particularly focusing on Bengali as a low-resource language and English as a high-resource benchmark. CrosGrpsABS constructs two complementary graphs: a syntactic graph that captures local grammatical relationships using dependency parsing tailored to the linguistic characteristics of Bengali, and a semantic graph that captures global contextual relationships by computing cosine similarities between transformer-based word embeddings. These graphs are dynamically fused through a bidirectional cross-attention mechanism that aligns and refines node representations based on cross-graph signals—allowing syntactic and semantic cues to reinforce each other. This design not only enhances sentiment prediction but also offers interpretability by highlighting influential syntactic structures and semantic associations.
@@
  • ARQ1: We present a unified graph construction strategy that captures both syntactic dependency and semantic similarity information using transformer-based contextual embeddings. These graphs are integrated through a novel bidirectional cross-attention mechanism, allowing the model to effectively leverage both syntactic structure and semantic context to improve the interpretability and performance of aspect-based sentiment analysis.
@@
  • ARQ2: We propose CrosGrpsABS, a hybrid ABSA framework that combines transformer-based and graph-based techniques, and demonstrate its adaptability to languages with varying resource availability. Through extensive experiments on a low-resource Bengali ABSA dataset and high-resource English datasets (SemEval 2014 Task 4: Restaurant and Laptop), we show that our model generalizes well across languages, highlighting its effectiveness in both constrained and rich-resource settings.
@@
The remainder of this paper is organized as follows. Section 2 reviews the related work on ABSA, graph-based models, and sentiment analysis for low-resource languages. Section 3 presents the architecture of the proposed CrosGrpsABS. Section 4 describes the experimental setup, including datasets and evaluation metrics. Section 5 discusses the results and compares them with existing baseline models. Finally, Section 6 concludes the study and suggests future research directions.
2 Literature Review
@@
Aspect-Based Sentiment Analysis in High-Resource Languages: Aspect-Based Sentiment Analysis (ABSA) has witnessed substantial progress in high-resource languages, particularly English, owing to the availability of large-scale annotated datasets, advanced language tools, and powerful pre-trained language models. Transformer-based architectures such as BERT [12], RoBERTa [13], and their specialized ABSA variants have significantly improved sentiment classification performance by capturing complex contextual dependencies through self-attention mechanisms [14]. These models often serve as the backbone for sentence-level and aspect-level classification tasks. Beyond sequential models, graph-based approaches have gained prominence for their ability to incorporate syntactic and semantic structures explicitly [7]. Dependency tree-based models [15, 16] and graph neural networks (GNNs) such as GCNs and TextGCNs [17, 18] have been employed to construct graph representations of sentence structure, enabling the learning of relational patterns between aspect and opinion terms. Semantic graphs based on similarity measures [19] further enhance global contextual understanding [20, 21]. However, many existing models in high-resource contexts treat syntactic and semantic features independently, limiting their capacity to leverage their complementary strengths for interpretability and performance.
Aspect-Based Sentiment Analysis in Low-Resource Languages (e.g., Bengali): Despite the global prevalence of languages like Bengali, ABSA research in low-resource settings remains underdeveloped. Bengali-specific sentiment analysis efforts have predominantly addressed coarse-grained tasks at the document or sentence level [6, 22], with limited exploration of fine-grained ABSA. The scarcity of annotated aspect-level datasets, lack of robust dependency parsers, and the unavailability of optimized Bengali-specific pre-trained models constrain the development of syntactic and semantic representations [23]. Furthermore, Bengali’s complex morphological structures and syntactic flexibility introduce challenges not observed in high-resource languages. Recent studies have attempted to adapt multilingual transformers (e.g., mBERT, XLM-R) for Bengali ABSA [24, 25], but performance often suffers due to mismatches in linguistic structure and domain specificity. Graph-based methods are even less explored in Bengali due to the difficulty of building accurate syntactic trees and semantic graphs without language-specific tools. This underscores the need for an adaptable ABSA framework that can overcome these limitations by integrating syntactic and semantic structures.
Research Gaps and Mitigation: Although both syntactic and semantic cues are vital for ABSA, most existing works treat them as disjoint components. Some recent research has attempted to bridge this gap using attention or fusion mechanisms, but few offer a unified and interpretable integration. Models that employ both syntactic dependencies and semantic similarities often fail to capture cross-level interactions in a structured and bidirectional manner [19, 26]. Furthermore, many of these models lack explainability, making it difficult to interpret how graph structures influence prediction outcomes. The problem becomes more pronounced in low-resource settings, where the challenges of representation construction are magnified. The absence of unified graph fusion strategies, especially for low-resource languages like Bengali, limits the development of generalized ABSA models. Moreover, cross-lingual adaptability has rarely been explored through graph-based techniques, which further restricts their application in multilingual contexts. Thus, a clear research gap exists in designing an interpretable graph-based ABSA framework that can jointly leverage syntactic structure and semantic context while maintaining adaptability across languages. This motivates the development of our proposed CrosGrpsABS model (i.e., Figure 2), which unifies syntactic (i.e., Section 3.1) and semantic (i.e., Section 3.2) graphs using a novel bidirectional cross-attention mechanism (i.e., Section 3.4, and Figure 4) to address both RQ1 (graph interaction shown in Figure 3 and interpretability shown in Figure 5) and RQ2 (cross-lingual adaptability), showing promising results (i.e., Section 5) in both Bengali (i.e., Table 5, 7, and 8) and English ABSA benchmarks (i.e., Table 6).
3 Proposed Methodology
@@
Aspect-based sentiment analysis (ABSA) aims to determine the sentiment polarity expressed toward specific aspects within a given sentence. Formally, consider a sentence $s=\{w_{1},w_{2},\ldots,w_{n}\}$ and an aspect $a=\{a_{1},a_{2},\ldots,a_{m}\}\subset s$. Here, $m$ denotes the end position of the aspect in $s$. When a sentence contains multiple aspects, each aspect is treated as a separate instance for sentiment classification. The goal is to predict the sentiment polarity (e.g., positive, negative, or neutral) associated with each aspect $a$. Figure 2 illustrates the overall architecture of CrosGrpsABS, which addresses the aforementioned problem setting. Given a sentence-aspect pair $(s,a)$, each token in the sentence is first mapped to its corresponding embedding through an embedding table $\mathbf{E}\in\mathbb{R}^{|V|\times d_{e}},$ where $|V|$ is the vocabulary size and $d_{e}$ is the embedding dimension. We concatenate the sentence and aspect to form a single input sequence for a transformer encoder (e.g., BERT [8] or XLM-RoBERTa [9]), formatting it as: $\text{Input}=[\text{[CLS]}]\;w_{1},\ldots,w_{n}\;[\text{SEP}]\;a_{1},\ldots,a_{m}\;[\text{SEP}],$ where [CLS] serves as a classification token and [SEP] tokens separate the sentence and aspect. Passing this sequence through the transformer yields contextual embeddings $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{T}]\in\mathbb{R}^{T\times d},$ with $T=n+m+3$ accounting for the added special tokens, and $d$ indicating the hidden dimensionality.
@@ 3.1 Syntactic Graph
3.2 Semantic Similarity Graph
@@
3.3 Graph Attention Networks (GAT)
@@
where $\mathbf{W}$ is a learnable weight matrix that projects node features into an attention space, and $\mathbf{a}$ is a learnable attention vector. The concatenation $\mathbf{W}\mathbf{h}_{i}\,\|\,\mathbf{W}\mathbf{h}_{j}]$ allows the attention mechanism to model interactions between the node and its neighbors, as shown in Equation (3). A LeakyReLU activation function is applied before computing the softmax normalization across all neighbors.
@@
By leveraging GAT, each token dynamically attends to its most relevant neighbors rather than treating all connected nodes equally. The attention mechanism in Equation (3) allows the model to assign higher weights to influential words, while Equation (4) ensures that only meaningful information is aggregated. Furthermore, stacking multiple GAT layers enables the model to capture multi-hop dependencies, which is essential for processing complex Bengali sentences. Ultimately, the combination of transformer-based self-attention with graph-based neighbor attention results in a richer and more structured representation of the input text.
3.4 Bidirectional Cross-Attention
@@
3.5 Transformer Encoder Refinement
@@
3.6 Aspect Representation Extraction
@@
where $|M|$ denotes the number of tokens forming the aspect, and $1,\ldots,M]$ are the corresponding indices in $\mathbf{H}_{\text{aspect}}$. This pooling step yields a compact embedding $\mathbf{z}_{\text{aspect}}$ that concentrates on the region of the sentence determining sentiment polarity. By focusing on aspect-specific dependencies, the model can disentangle general context from cues most pertinent to the aspect, ultimately enhancing its ability to capture fine-grained sentiment relationships in ABSA.
Table 1: Data distribution for the four domain-specific ABSA datasets in Bangla.
@@
Table 2: Statistics of SemEval 2014 Task 4 datasets (Laptop and Restaurant) with sentiment distribution.
@@
Table 3: Hyperparameter Settings and Additional Techniques
@@
Table 4: Comparative performance of three embedding-based approaches (mBERT, RoBERTa, BanglaBERT) on four domain-specific Bengali ABSA datasets. Acc and F1 are reported in percentage. The best results in each column are highlighted in bold.
@@
$uparrow Acc\uparrow F1\uparrow$ $uparrow Acc\uparrow F1\uparrow$
@@
Table 5: Performance comparison of our proposed model with baseline models on multiple datasets. The best results are highlighted in bold, and the second-best results are underlined.
@@
$uparrow Acc\uparrow F1\uparrow$ $uparrow Acc\uparrow F1\uparrow$
Roberta-base (Zero-shot) 50.75 46.51 52.53 48.59 63.24 59.07 57.65 51.68
Roberta-large (Zero-shot) 54.52 57.91 68.82 67.46 66.08 65.58 71.82 70.68
Roberta 56.34 50.41 68.28 68.16 74.32 74.46 80.13 79.91
BanglaBERT 56.81 41.16 48.39 31.56 68.24 55.36 54.97 38.99
@@
3.7 Highway Gating Mechanism
$\bar{\mathbf{h}}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{H}_{\text{refined}}[i],$ (14)
Table 6: Comparison of Different Models on ABSA Tasks (SemEval 2014 Task 4 datasets: Restaurant and Laptop)
@@ 3.7 Highway Gating Mechanism
Table 7: Ablation study results for our proposed model across multiple datasets. The best results are highlighted in bold.
@@
We then define two functions acting on $\mathbf{u}$. The first, $\mathbf{H}(\cdot)$, is a feed-forward transformation (e.g., a ReLU-activated linear layer) that generates a candidate set of new features. The second, $\mathbf{T}(\cdot)$, is a gating function that produces a sigmoid output indicating how much of these new features should be integrated into the final representation [39]. Specifically, we compute
$\mathbf{z}_{H}=\mathbf{T}(\mathbf{u})\,\odot\,\mathbf{H}(\mathbf{u})\;+\;\bigl(\mathbf{1}-\mathbf{T}(\mathbf{u})\bigr)\,\odot\,\mathbf{u},$ (16)
where $\mathbf{T}(\mathbf{u})=\sigma\bigl(\mathbf{W}_{T}\mathbf{u}+\mathbf{b}_{T}\bigr)$. In Equation (16), the gating vector $\mathbf{T}(\mathbf{u})$ determines the balance between newly transformed information $\mathbf{H}(\mathbf{u})$ and the original input $\mathbf{u}$. As a result, tokens closely related to the aspect can be emphasized, while broader contextual cues remain accessible. This design also helps mitigate vanishing gradients by allowing unimpeded flow of input features (i.e., skip connections), making the gating mechanism robust in deeper networks [33, 39]. Consequently, the final output $\mathbf{z}_{H}$ represents a balanced fusion of aspect-specific cues and the preserved global context, thereby enhancing the model’s ability to capture nuanced sentiment relationships.
3.8 Classification Layer